A circus act consists of a performer (a "human cannonball") being launched from a cannon into a net at the other end of the stage. The performer (mass 60 kg) leaves the cannon at a speed of 25 m/s at an angle of 30o above horizontal. She comes out of the cannon at a height of 2 meters above the ground and lands in the net at a horizontal distance of 45 meters away from the launch point.

1. How long is the performer in the air before hitting the net? Answer in seconds.

2. How high above the ground is the net? Answer in meters.

3. What is the x component of the performer's velocity just before she hits the net? Answer in m/s.

4. What is the human cannonball's speed just before she hits the net? Answer in m/s.

Hint: the speed is the magnitude of the velocity. You will need to find the y component of the velocity.
Answer #1

1. sx = xt {as ax=0] 45M = (25ml8 cas10) t & t= 2.088 Y-Y. = Viyt & blyt h-2= (258ingo) (2.08) + Th= 6.78m (-9-8)(2:083 3. =